# Automatically Detecting Online Deceptive Patterns

### v3

@@ 4.2 Web-UI Element Detection
[…]
@@
While Web-UI Element Detection is a well-studied problem, the existing approaches face several limitations in our deceptive pattern detection task. Detectors like Ominparser 2 [44], Ferret UI 2 [42], UEID [80], and Element Detector [79] do not distinguish between checked and unchecked states. Detectors like UISketch [62] report low performance on real-world websites. ScreenRecognition [84] shows promise for this step but was trained on mobile screenshots and is closed-source, preventing its use or testing on websites.
@@
To improve the detector performance, we explored modifying existing datasets used to train them. Omniparser 2 uses a manually annotated dataset of popular websites, but the dataset is not public [44]. UEID [80] was trained on the RICO Dataset [43], which contains manually annotated mobile app UIs. We could not modify this dataset to train a detector for webpages. Ferret UI 2 [42] and Element Detector [79] use WebUI [79], a popular framework with 400k websites which is publicly available, but the automatically computed labels from accessibility trees [^4] Accessibility trees are generated using aria labels which developers have to optionally add to a website. are noisy and miss a lot of elements on a given website [79, 78]. UISketch contains hand-drawn images to identify Web-UI Elements based on their sketches and does not generalize well to real-world websites [62].
@@ 4.2.1 Dataset Curation
We develop a novel approach to create a diverse and representative dataset to address the lack of suitable Web-UI Element Dataset for training our Web-UI Element Detector, as shown in Figure 4. Instead of relying on manual scraping and labeling [84, 62, 81, 44], we leverage AI-powered tools to generate a synthetic dataset that reflects the current web landscape. Our approach allows for precise control over element positioning and labeling, which allows us to scale the dataset generation.
We generated 2.5K ideas for diverse websites using GPT-4 [5]. We passed these ideas to v0555https://v0.dev/ (refer to Section 11.4 for examples), an AI-based website generator, to generate three websites per idea. v0 uses shadcn [63], a customizable UI-Library, which allows us to get bounding boxes and the state (checked/unchecked) of every UI element. Three members of the research team manually verified the 7.5K websites generated by v0. With a failure rate of less than 2%, this process was relatively fast, as errors in websites are detected by the compiler. These errors consisted mainly of typos and missing libraries that the researchers manually fixed. We then rendered each v0 generated website (7.5K) using randomized components from 6 popular UI-libraries, like Material UI and Bootstrap [^6] https://mui.com/material-ui/, https://getbootstrap.com/. This process resulted in 62K screenshots, where each screenshot had a bounding box and a label for each UI element. Recall that we have 7 labels for the UI elements. We refer to this dataset as the Web-UI Elements Dataset.
@@ 4.2.2 Training Web-UI Element Detector
We adopt the YOLOv10 model (You Only Look Once (YOLO) architecture [75], a real-time object detector for recognizing UI elements from a screenshot. We chose YOLOv10 over Convolution Neural Networks (CNNs) [35, 46, 18] and VLLMs like Molmo [25]. CNN-based detectors, such as Faster R-CNN [57], provide predictions with high accuracy, but require considerable computational power and time [56, 57]. VLLMs, despite their strong capabilities in understanding image context, demonstrate significant limitations in image classification tasks [85], specifically object detection tasks [83]. Evaluating Molmo [25] on detecting Web-UI Elements yielded poor results, as shown in Table 1. YOLO models are comparatively lightweight, around 40MB, allowing for various deployment options without requiring extensive compute resources.
@@
[…]
@@ 4.4.1 Vision Dataset
[…]
@@ 4.4.2 Vision Module Evaluation
Table 1: Performance of the YOLOv10 Ensemble and Molmo [25] on our Real-UI Dataset
@@
Table 1 shows the F1-scores of our YOLOv10 Ensemble and Molmo [25] for each of the 7 classes. We observe that the ensemble outperforms Molmo for all the UI elements. Note that few-shot prompting of local VLLMs is not possible for image inputs. The ensemble exhibits a relatively lower performance on the unchecked radio button and unchecked check box, because they look visually similar to the other UI elements.
@@ 5.1 Possible Solutions
Prior works have shown that LLMs [49] are suitable for reasoning tasks similar to our task. These models fall into three categories: 1) large Language Models (LMs) like Gemini and GPT-4, 2) small LMs such as Gemma and Qwen, and 3) very small LMs like T5. These models have varying capabilities and trade-offs, as described in Table 2.
@@ 5.1.1 Large Language Models (LLMs):
LLMs are very effective at performing a wide range of tasks [23, 7, 66] and are able to closely follow user instructions [52, 86]. However, as shown in Table 2, using these LLMs is cost-prohibitive and has potential data privacy concerns [3].
@@ 5.1.2 Small Language Models (SLMs):
Unlike LLMs that follow complex instructions when performing a task, SLMs often struggle to do so. This limitation of SLMs can be overcome by fine-tuning them on specific tasks, improving their performance to match that of LLMs [45]. Moreover, as shown in Table 2, SLMs have two key advantages over LLMs: 1) being compute efficient, they can be deployed locally across a wide range of platforms, and 2) they alleviate any data privacy concern associated with API based LLMs.
@@ 5.1.3 Very Small Language Models (vSLMs):
SLMs, while compute-efficient, are not the best solution to use in an extremely resource-constrained platform, such as in-browser or on devices with no specialized GPU. In such environments, we can leverage vSLMs like Flan-T5 [20]. Models like Flan-T5 need to be finetuned or distilled from LLMs for specific tasks to achieve high accuracy [33]. We show the detailed distillation steps we performed in Section 5.2.4.
In summary, we observe that large LMs, such as Gemini, are considerably effective in detecting deceptive patterns from an ElementMap. However, as described in Table 2, utilizing such large and closed-source models presents challenges like high usage cost, considerable latency, and potential data-privacy concerns as the ElementMap is sent to an external service. Smaller LMs such as Qwen and T5 address these challenges and have been shown to perform well on such specific tasks after finetuning [45, 33].
@@ 5.2.1 Prompting the LLM
We utilize proven techniques like Chain-of-Thought (CoT) [77], few-shot prompting [15], and prompting the model to reason about its classification [51, 32, 49] to help the model understand the task and identify deceptive patterns. Specifically, our system prompt, $P_{\text{system}}$ [^7] Please find the system prompt under Files here: https://osf.io/tha2d/?view_only=4fd2116fa2e94c99857679eddfea5937, provides the LLM with a plan on how to detect deceptive patterns and instructs the LLM to generate three columns — Deceptive Category, Deceptive Subtype, and Reasoning — for each element in the ElementMap, as shown in Figure 5. This prompt allows us to not only generate precise labels and the associated reasoning but also minimize hallucinations. An added benefit of generating the classification with ‘Reasoning’ is that we can use these ‘rationales’ to further train smaller LMs. This Distilling Step-by-Step methodology has been shown to be very effective by Hsieh et al. [33].
During our initial evaluation of Gemini 1.5 Pro, using $P_{\text{system}}$, we observed that the model could identify all deceptive patterns in the ElementMap, but would often mis-classify ‘non-deceptive’ patterns as deceptive, resulting in a high false positive rate and a low precision score. To address this problem, we utilized the Gemini 2.0-Flash-Thinking reasoning model. Reasoning models have shown promising results in being able to reason about a task [45, 51]. We prompted the LLM to re-evaluate the elements identified as deceptive and correct misclassifications7. Next, we utilized the Gemini 2.0-Flash-Thinking model to re-verify the labels on every website with one or more elements classified as deceptive. This additional step in generating the labels, significantly reduced the number of false positives we observed. We note that we do not use this model as our base model because of its limited availability, making it infeasible to be used at scale. We show that this approach results in high precision and recall in Section 5.3.
@@ 5.2.2 Creating Distillation Dataset
We create $\mathcal{D_{\text{distill}}}$ by scraping and analyzing 11K websites from the Tranco list [54]. We filter out non-English websites using the langdetect library [67] and adult websites using the NudeNet package [50], leaving us with 6,626 websites. For these websites, a significant portion was non-deceptive or had very few deceptive patterns. To increase websites with deceptive patterns, we opted to look at e-commerce websites, as these websites tend to have deceptive patterns [47]. As such, we analyze an additional 4,492 featured websites from Shopify Partners Directory [^8] https://www.shopify.com/partners/directory.
Overall, we have 11,118 websites in our dataset. We run AutoBot on these websites, using the Gemini 1.5 Pro and Gemini 2.0 Flash-Thinking models. To reduce randomness and have deterministic outputs, we limit the $temperature=0$ and $top_{p}=0.1$ [58]. We incorporate the final classification and reasoning produced by AutoBot in $\mathcal{D_{\text{distill}}}$. The distribution of samples (a sample is defined as a single row of the ElementMap) is shown in Table 3.
@@
DeepScaleR [45] has shown that small language models (SLMs) finetuned for specific tasks are able to mirror the performance of larger models on those same tasks. Using this insight, we distill a Qwen2.5-1.5B model to be able to mimic Gemini’s performance at detecting deceptive patterns. We use the $\mathcal{D_{\text{distill}}}$ dataset to perform full-finetuning of the Qwen2.5-1.5B model. As shown in Table 3, the distribution of samples in the distillation dataset is highly imbalanced, with over 92% of the samples being ‘non-deceptive’. Training on such an imbalanced dataset may introduce bias towards the majority class [38]. To mitigate such bias, we perform Random Under Sampling of the ‘non-deceptive’ class, which has shown to improve performance [26], to achieve a more balanced distribution of about 55% ‘non-deceptive’ samples. Our distilled version of Qwen2.5-1.5B was trained on 34.7M tokens on 4 Nvidia-A6000 GPUs for 2 epochs. We present the performance of the trained SLM in Section 5.3.
@@ 5.2.4 Distilling Very Small Language Models (T5):
To distill the knowledge from LLMs into very small language models like T5, we use and improve the paradigm introduced by Hsieh et al. [33]. We split $\mathcal{D_{\text{distill}}}$ into 90% training and a 10% validation set grouped by the sites from which the samples were extracted. This ensures that samples from the same site are not present in training and testing sets. We formally define the training dataset, $D_{\text{train}}$, as:
@@ 6.1 Dataset
To create the dataset for the end-to-end evaluation, we crawl the websites in the deceptive pattern dataset from Mathur et al. [47]. We use this dataset since it is the latest and most comprehensive dataset of websites with deceptive patterns. We filter out the non-English and offline websites from this dataset, leaving us with 555 websites out of the 1400 sites mentioned. Next, we crawled the top 1000 websites from the Tranco [54] list [^9] https://tranco-list.eu/list/QGJ74/1000000, utilizing the same methodology in Section 5.2.2 to filter out non-English and NFSW websites, resulting in 597 websites. Additionally, for the Tranco websites, we did not scrape just the landing pages of websites. Rather, we queried “site:domain” on DuckDuckGo and programmatically counted the number of interactable elements for each of the top ten resulting pages. The page with the highest count was selected for analysis. In total, we have 1152 websites in our end-to-end evaluation dataset.
Next, we manually labeled the screenshots of these websites to identify and localize the deceptive patterns, associating each screenshot with an ElementMap. Specifically, two authors annotated the same randomly selected 100 websites with high inter-annotator agreement ($\kappa=0.89$[37]. Then, each researcher separately labeled the rest of the screenshots. We refer to this dataset as the golden Deceptive Design Dataset (D3 dataset). We show the distribution of the samples in D3 Dataset in Table 5.
@@
Table 6: Performance of AutoBot (with three underlying language models: Gemini, Qwen, and T5) and DPGuard [64] on the D3 Dataset. The table has three sections, each showing performance at the category, subtype, and binary levels.
@@
Gemini Qwen T5 DPGuard [64] Gemini Qwen T5 DPGuard [64] Gemini Qwen T5 DPGuard [64]

### v4

@@ 4.2 Web-UI Element Detection
The Web-UI Element Detection step uses the same screenshot to identify and localize these 7 Web-UI Elements: buttons, checkboxes (, ), radio buttons (ȷ, ), and toggle switches (v, u). This step provides context to the extracted text and enables a more comprehensive understanding of the webpage’s structure. For instance, distinguishing between a clickable button and a static text block can be significant, as a seemingly simple text might be a deceptive call to action when recognized as a button. Similarly, identified checkbox states (checked or unchecked) can reveal pre-selected options that users might overlook.
@@
While Web-UI Element Detection is a well-studied problem, the existing approaches face several limitations in our deceptive pattern detection task. Detectors like Ominparser 2 [lu2024omniparser], Ferret UI 2 [li2024ferret], UEID [xie2020uied], and Element Detector [wu2023webui] do not distinguish between checked and unchecked states. Detectors like UISketch [sermuga2021uisketch] report low performance on real-world websites. ScreenRecognition [zhang2021screen] shows promise for this step but was trained on mobile screenshots and is closed-source, preventing its use or testing on websites.
@@
To improve the detector performance, we explored modifying existing datasets used to train them. Omniparser 2 uses a manually annotated dataset of popular websites, but the dataset is not public [lu2024omniparser]. UEID [xie2020uied] was trained on the RICO Dataset [rico], which contains manually annotated mobile app UIs. We could not modify this dataset to train a detector for webpages. Ferret UI 2 [li2024ferret] and Element Detector [wu2023webui] use WebUI [wu2023webui], a popular framework with 400k websites which is publicly available, but the automatically computed labels from accessibility trees [^4] Accessibility trees are generated using aria labels which developers have to optionally add to a website. are noisy and miss a lot of elements on a given website [wu2023webui, wu_never-ending_2023]. UISketch contains hand-drawn images to identify Web-UI Elements based on their sketches and does not generalize well to real-world websites [sermuga2021uisketch].
@@ 4.2.1 Dataset Curation
We develop a novel approach to create a diverse and representative dataset to address the lack of suitable Web-UI Element Dataset for training our Web-UI Element Detector, as shown in Figure 4. Instead of relying on manual scraping and labeling [zhang2021screen, sermuga2021uisketch, ueid_xie, lu2024omniparser], we leverage AI-powered tools to generate a synthetic dataset that reflects the current web landscape. Our approach allows for precise control over element positioning and labeling, which allows us to scale the dataset generation.
We generated 2.5K ideas for diverse websites using GPT-4 [achiam2023gpt]. We passed these ideas to v0555https://v0.dev/ (refer to Section 11.4 for examples), an AI-based website generator, to generate three websites per idea. v0 uses shadcn [shadcn], a customizable UI-Library, which allows us to get bounding boxes and the state (checked/unchecked) of every UI element. Three members of the research team manually verified the 7.5K websites generated by v0. With a failure rate of less than 2%, this process was relatively fast, as errors in websites are detected by the compiler. These errors consisted mainly of typos and missing libraries that the researchers manually fixed. We then rendered each v0 generated website (7.5K) using randomized components from 6 popular UI-libraries, like Material UI and Bootstrap [^6] https://mui.com/material-ui/, https://getbootstrap.com/. This process resulted in 62K screenshots, where each screenshot had a bounding box and a label for each UI element. Recall that we have 7 labels for the UI elements. We refer to this dataset as the Web-UI Elements Dataset.
@@ 4.2.2 Training Web-UI Element Detector
We adopt the YOLOv10 model (You Only Look Once (YOLO) architecture [wang2024yolov10], a real-time object detector for recognizing UI elements from a screenshot. We chose YOLOv10 over Convolution Neural Networks (CNNs) [khandelwal_prisec_2021, mansur2023aidui, chen2023unveiling] and VLLMs like Molmo [molmo]. CNN-based detectors, such as Faster R-CNN [ren2015faster], provide predictions with high accuracy, but require considerable computational power and time [redmon2018yolov3, ren2015faster]. VLLMs, despite their strong capabilities in understanding image context, demonstrate significant limitations in image classification tasks [zhang2024visually], specifically object detection tasks [zang2025contextual]. Evaluating Molmo [molmo] on detecting Web-UI Elements yielded poor results, as shown in Table 1. YOLO models are comparatively lightweight, around 40MB, allowing for various deployment options without requiring extensive compute resources.
@@
During training, we observed that YOLOv10 models could not distinguish between the 7 labels (first column of Table 1) accurately. We attribute the reason to the labels being visually similar, such as a checked switch and a checked radio button. As such, we trained three YOLOv10 models, each focused on distinguishing between 2-3 different elements. The first model labeled button and ȷ; the second labeled v, , and ; and the third labeled u and . We found that each model performed much better than one trying to handle all the classes at once. The same observation has been made in literature before for YOLO-based object detection [mohankumar_benchmark_nodate, walambe_lightweight_2021, pham_optimizing_2024]. We combined the outputs of the three models by simply performing a union over the detected UI elements. In case of overlap, we took the label with the higher confidence classification. We refer to this detection method as the YOLOv10 Ensemble.
@@ 4.4.1 Vision Dataset
We curate a labeled dataset of UI elements from the deceptive pattern websites dataset of Mathur et al. [mathur2019dark]. We manually annotated over 1.5K website screenshots using Label Studio [Label_Studio]. In particular, one author manually annotated each screenshot by drawing bounding boxes around each UI element and assigning it a type. The type is one the 7 Web-UI Elements: buttons, checkboxes (, ), radio buttons (ȷ, ), and toggle switches (v, u). Another author independently verified the annotations. Both authors then discussed and resolved the conflicts in annotations. We refer to this dataset as the Real-UI Dataset.
@@ 4.4.2 Vision Module Evaluation
Table 1: Performance of the YOLOv10 Ensemble and Molmo [molmo] on our Real-UI Dataset
@@
Table 1 shows the F1-scores of our YOLOv10 Ensemble and Molmo [molmo] for each of the 7 classes. We observe that the ensemble outperforms Molmo for all the UI elements. Note that few-shot prompting of local VLLMs is not possible for image inputs. The ensemble exhibits a relatively lower performance on the unchecked radio button and unchecked check box, because they look visually similar to the other UI elements.
@@ 5.1 Possible Solutions
Prior works have shown that LLMs [nayak2024experimental] are suitable for reasoning tasks similar to our task. These models fall into three categories: 1) large Language Models (LMs) like Gemini and GPT-4, 2) small LMs such as Gemma and Qwen, and 3) very small LMs like T5. These models have varying capabilities and trade-offs, as described in Table 2.
@@ 5.1.1 Large Language Models (LLMs):
LLMs are very effective at performing a wide range of tasks [cui2025curie, agarwal2024llm, shool2025systematic] and are able to closely follow user instructions [ouyang2022training, zhou2023instruction]. However, as shown in Table 2, using these LLMs is cost-prohibitive and has potential data privacy concerns [api_security].
@@ 5.1.2 Small Language Models (SLMs):
Unlike LLMs that follow complex instructions when performing a task, SLMs often struggle to do so. This limitation of SLMs can be overcome by fine-tuning them on specific tasks, improving their performance to match that of LLMs [deepscaler2025]. Moreover, as shown in Table 2, SLMs have two key advantages over LLMs: 1) being compute efficient, they can be deployed locally across a wide range of platforms, and 2) they alleviate any data privacy concern associated with API based LLMs.
@@ 5.1.3 Very Small Language Models (vSLMs):
SLMs, while compute-efficient, are not the best solution to use in an extremely resource-constrained platform, such as in-browser or on devices with no specialized GPU. In such environments, we can leverage vSLMs like Flan-T5 [FlanT5]. Models like Flan-T5 need to be finetuned or distilled from LLMs for specific tasks to achieve high accuracy [hsieh2023distilling]. We show the detailed distillation steps we performed in Section 5.2.4.
In summary, we observe that large LMs, such as Gemini, are considerably effective in detecting deceptive patterns from an ElementMap. However, as described in Table 2, utilizing such large and closed-source models presents challenges like high usage cost, considerable latency, and potential data-privacy concerns as the ElementMap is sent to an external service. Smaller LMs such as Qwen and T5 address these challenges and have been shown to perform well on such specific tasks after finetuning [deepscaler2025, hsieh2023distilling].
@@ 5.2.1 Prompting the LLM
We utilize proven techniques like Chain-of-Thought (CoT) [wei2022chain], few-shot prompting [brown2020language], and prompting the model to reason about its classification [openai_learning_reason_2024, guo2025deepseek, nayak2024experimental] to help the model understand the task and identify deceptive patterns. Specifically, our system prompt, $P_{\text{system}}$ [^7] Please find the system prompt under Files here: https://osf.io/tha2d/?view_only=4fd2116fa2e94c99857679eddfea5937, provides the LLM with a plan on how to detect deceptive patterns and instructs the LLM to generate three columns — Deceptive Category, Deceptive Subtype, and Reasoning — for each element in the ElementMap, as shown in Figure 5. This prompt allows us to not only generate precise labels and the associated reasoning but also minimize hallucinations. An added benefit of generating the classification with ‘Reasoning’ is that we can use these ‘rationales’ to further train smaller LMs. This Distilling Step-by-Step methodology has been shown to be very effective by Hsieh et al. [hsieh2023distilling].
During our initial evaluation of Gemini 1.5 Pro, using $P_{\text{system}}$, we observed that the model could identify all deceptive patterns in the ElementMap, but would often mis-classify ‘non-deceptive’ patterns as deceptive, resulting in a high false positive rate and a low precision score. To address this problem, we utilized the Gemini 2.0-Flash-Thinking reasoning model. Reasoning models have shown promising results in being able to reason about a task [deepscaler2025, openai_learning_reason_2024]. We prompted the LLM to re-evaluate the elements identified as deceptive and correct misclassifications7. Next, we utilized the Gemini 2.0-Flash-Thinking model to re-verify the labels on every website with one or more elements classified as deceptive. This additional step in generating the labels, significantly reduced the number of false positives we observed. We note that we do not use this model as our base model because of its limited availability, making it infeasible to be used at scale. We show that this approach results in high precision and recall in Section 5.3.
@@ 5.2.2 Creating Distillation Dataset
We create $\mathcal{D_{\text{distill}}}$ by scraping and analyzing 11K websites from the Tranco list [pochat2018tranco]. We filter out non-English websites using the langdetect library [langdetect] and adult websites using the NudeNet package [nudenet], leaving us with 6,626 websites. For these websites, a significant portion was non-deceptive or had very few deceptive patterns. To increase websites with deceptive patterns, we opted to look at e-commerce websites, as these websites tend to have deceptive patterns [mathur2019dark]. As such, we analyze an additional 4,492 featured websites from Shopify Partners Directory [^8] https://www.shopify.com/partners/directory.
Overall, we have 11,118 websites in our dataset. We run AutoBot on these websites, using the Gemini 1.5 Pro and Gemini 2.0 Flash-Thinking models. To reduce randomness and have deterministic outputs, we limit the $temperature=0$ and $top_{p}=0.1$ [renze2024effect]. We incorporate the final classification and reasoning produced by AutoBot in $\mathcal{D_{\text{distill}}}$. The distribution of samples (a sample is defined as a single row of the ElementMap) is shown in Table 3.
@@
DeepScaleR [deepscaler2025] has shown that small language models (SLMs) finetuned for specific tasks are able to mirror the performance of larger models on those same tasks. Using this insight, we distill a Qwen2.5-1.5B model to be able to mimic Gemini’s performance at detecting deceptive patterns. We use the $\mathcal{D_{\text{distill}}}$ dataset to perform full-finetuning of the Qwen2.5-1.5B model. As shown in Table 3, the distribution of samples in the distillation dataset is highly imbalanced, with over 92% of the samples being ‘non-deceptive’. Training on such an imbalanced dataset may introduce bias towards the majority class [leevy2018survey]. To mitigate such bias, we perform Random Under Sampling of the ‘non-deceptive’ class, which has shown to improve performance [elsoud2024under], to achieve a more balanced distribution of about 55% ‘non-deceptive’ samples. Our distilled version of Qwen2.5-1.5B was trained on 34.7M tokens on 4 Nvidia-A6000 GPUs for 2 epochs. We present the performance of the trained SLM in Section 5.3.
@@ 5.2.4 Distilling Very Small Language Models (T5):
To distill the knowledge from LLMs into very small language models like T5, we use and improve the paradigm introduced by Hsieh et al. [hsieh2023distilling]. We split $\mathcal{D_{\text{distill}}}$ into 90% training and a 10% validation set grouped by the sites from which the samples were extracted. This ensures that samples from the same site are not present in training and testing sets. We formally define the training dataset, $D_{\text{train}}$, as:
@@ 6.1 Dataset
To create the dataset for the end-to-end evaluation, we crawl the websites in the deceptive pattern dataset from Mathur et al. [mathur2019dark]. We use this dataset since it is the latest and most comprehensive dataset of websites with deceptive patterns. We filter out the non-English and offline websites from this dataset, leaving us with 555 websites out of the 1400 sites mentioned. Next, we crawled the top 1000 websites from the Tranco [pochat2018tranco] list [^9] https://tranco-list.eu/list/QGJ74/1000000, utilizing the same methodology in Section 5.2.2 to filter out non-English and NFSW websites, resulting in 597 websites. Additionally, for the Tranco websites, we did not scrape just the landing pages of websites. Rather, we queried “site:domain” on DuckDuckGo and programmatically counted the number of interactable elements for each of the top ten resulting pages. The page with the highest count was selected for analysis. In total, we have 1152 websites in our end-to-end evaluation dataset.
Next, we manually labeled the screenshots of these websites to identify and localize the deceptive patterns, associating each screenshot with an ElementMap. Specifically, two authors annotated the same randomly selected 100 websites with high inter-annotator agreement ($\kappa=0.89$) [landis1977measurement]. Then, each researcher separately labeled the rest of the screenshots. We refer to this dataset as the golden Deceptive Design Dataset (D3 dataset). We show the distribution of the samples in D3 Dataset in Table 5.
@@
Table 6: Performance of AutoBot (with three underlying language models: Gemini, Qwen, and T5) and DPGuard [50Shades] on the D3 Dataset. The table has three sections, each showing performance at the category, subtype, and binary levels.
@@
Gemini Qwen T5 DPGuard [50Shades] Gemini Qwen T5 DPGuard [50Shades] Gemini Qwen T5 DPGuard [50Shades]